Grade 11Physical Chemistry

A mixture containing 100 gm H2 and 100 gm O2 is ignited so that water is formed according to the reaction, 2H2 + O2 - 2H2O; How much water will be formed

khushbu Sharma
9 Years agoGrade 11

1 Answer

Nayana
9 Years ago
Moles of H2=100÷2=50..moles of oxygen=100÷32=3.125..the limiting reagent is clearly O2.therefore according to the equation ...1 mole of O2 gives 2 mole of water.so 3.125 moles of O2 gives =2×3.125 moles of H20. Moles of H20=6.25.amount of H20=18×6.25=112.5g